OpenStudy (anonymous):

Estimate by using compatible number 105(2/3)/26(1/2)=___________ The () are the fraction part of the number I just didnt know how to put it in right.

OpenStudy (anonymous):

105(2/3) -------- 26(1/2) = 35(2) ------ 13 = 70/13